OpenStudy (anonymous):

check my answer? Find S_11 for 1 + 2 + 4 + 8 + ... I got 2047, is that correct?

OpenStudy (anonymous):

wht is "S"?

OpenStudy (anonymous):

I believe that it means the sum of the series

OpenStudy (anonymous):

idk the answer to this

OpenStudy (sloppycanada):

Write out your equation for me.

OpenStudy (anonymous):

well I know that r=2 and so I multiplied by 2 for each and got 1,2,4,8,16,32,64,128,256,512,1024 and then I looked it up because I wasn't sure what to do next and they got 2047 and I wanted to know how they got that

OpenStudy (sloppycanada):

There is an equation. First you have to find the t11. T11 = term 11.

OpenStudy (anonymous):

The eleventh term is 1024 I believe

OpenStudy (anonymous):

its geometric progression, a, ar, ar^2, ar^3.... the common ratio here is 2, that is, every next term is "previous term multiplied by 2" sum in a GP is given by the formula \[ \frac{ a (1-r ^{n})}{ 1-r }\]

OpenStudy (anonymous):

take a as 1, r as 2 and n as 11

OpenStudy (anonymous):

(1(1-2^11))/1-2 =1-2048/1-2 =2047

OpenStudy (sloppycanada):

So yes, your answer is 2047

OpenStudy (anonymous):

okay thank you